OpenStudy (bbcream14):

I will fan Solve the following system of equations by using elimination 1.) -x+y=7 X+y=1 2.) 5x-2y=24 3x+2y=24

OpenStudy (bbcream14):

Fan and metal

Vocaloid (vocaloid):

let's start with the first one -x+y=7 x+y=1 let's try adding these equations together! what is -x + x?

OpenStudy (bbcream14):

0

Vocaloid (vocaloid):

good! what is y + y?

OpenStudy (bbcream14):

2

Vocaloid (vocaloid):

not quite! hint: A + A = 2A y + y = ? follow the pattern

OpenStudy (bbcream14):

So 2y

Vocaloid (vocaloid):

great! next step is to add 7 + 1 = 8 therefore: 2y = 8 what does y equal?

OpenStudy (bbcream14):

8

Vocaloid (vocaloid):

not quite! we know that "2y" is equal to 8. but what about "y" by itself?

OpenStudy (bbcream14):

4

Vocaloid (vocaloid):

good! so y = 4 now, we can go back to our original equations if x + y = 1, AND y = 4, then what is x?

OpenStudy (bbcream14):

1

Vocaloid (vocaloid):

not quite! x + y = 1 therefore x + 4 = 1 __ + 4 = 1 fill in the blank

OpenStudy (bbcream14):

-5

Vocaloid (vocaloid):

getting warmer but not quite

Vocaloid (vocaloid):

hint: subtract 4 from both sides 1 - 4 = ?

OpenStudy (bbcream14):

3

Vocaloid (vocaloid):

not quite...

Vocaloid (vocaloid):

use a calculator if you are unsure

OpenStudy (bbcream14):

0

Vocaloid (vocaloid):

nope, try again

Vocaloid (vocaloid):

use this calculator: https://www.google.co.uk/?gfe_rd=cr&ei=3WBAWN_0B-zR8gfOxIOoAg#q=calc

Vocaloid (vocaloid):

type in "1 - 4"

OpenStudy (bbcream14):

-3

Vocaloid (vocaloid):

good job! so our answer for #1 is (-3,4) lets move on to problem 2!

Vocaloid (vocaloid):

5x-2y=24 3x+2y=24 again, we can add these together 3x + 5x = ?

OpenStudy (bbcream14):

8x

Vocaloid (vocaloid):

good job! now what is -2y + 2y?

OpenStudy (bbcream14):

0

Vocaloid (vocaloid):

great job! next step is to add 24 + 24 = 48 now we have 8x = 48 what does "x" equal?

OpenStudy (bbcream14):

28

Vocaloid (vocaloid):

8 * ____ = 48 fill in the blank

OpenStudy (bbcream14):

6

Vocaloid (vocaloid):

good, so x = 6 next step: we know that 3x+2y=24 if we substitute x into the equation we get 18 + 2y = 24 so, we can subtract 18 from both sides to get 2y = 6 what does "y" equal?

OpenStudy (bbcream14):

3

Vocaloid (vocaloid):

good so the answer to #2 is (6,3)